# Gas price shocks, Uncertainty and Price setting: Evidences from Italian Firms

### v1

@@ Abstract
This paper examines how natural gas price shocks affect Italian firms’ pricing decisions and inflation expectations using quarterly survey data from the Bank of Italy’s Survey on Inflation and Growth Expectations (SIGE) spanning 1999Q4–2025Q2. We identify natural gas price shocks through a Bayesian VAR with sign and zero restrictions. Our findings reveal that these shocks are a primary driver of firms’ inflation expectations, particularly during the post-COVID period (2021–2023) when supply disruptions following Russia’s invasion of Ukraine generated unprecedented price pressures. We then estimate a larger BVAR incorporating firm-level price setting variables and macro aggregates, documenting that gas price shocks generate persistent increases in both firms’ current and expected prices, alongside elevated inflation uncertainty. We uncover substantial non-linearities using state-dependent local projections: under high uncertainty, firms successfully pass through cost increases to consumers, maintaining elevated prices; under low uncertainty, recessionary effects dominate, causing firms to reduce prices below baseline.
Keywords: Gas price shocks, Firms’ Price Setting, Inflation Uncertainty, Firms Expectations, Local Projections, Bayesian Vector Autoregression, Sign Restrictions
@@ 1 Introduction
The period that follows the Covid-19 pandemic has seen an unprecedented rise in inflation in European economies. The years that followed the post pandemic recovery have been characterized by strong political tensions, dominated by events such as Russian invasion of Ukraine in the first part of 2022 which led to massive disruptions in natural gas supplies to European countries, raising the attention on inflationary role of supply shocks affecting this important commodity. Due to the lack of domestic energy sources, Italy is a major importer and deeply reliant on natural gas as one of its main power sources, making it particularly sensitive to supply disruptions that followed this dramatic geopolitical turning point. Italian firms have been particularly exposed to the unprecedented increase in the price of natural gas, which had a substantial impact on the price of electricity in the country. We show that, unsurprisingly, this has led to an unprecedented increase in firms’ uncertainty about the future level of prices.
Unlike the extensive literature on oil supply shocks, research examining the role of natural gas supply disruptions has emerged only recently. In this paper, we aim at identifying a shock to the nominal price of natural gas and at evaluating the impact of this shock on firms price setting behavior. We identify the shock by adopting the behavioral Bayesian VAR model proposed by Kilian and Zhou (2022), on which we impose a combination of sign and zero restrictions. Then, we evaluate the effect of this shock on firms’ pricing choices using generalized impulse response functions from a larger BVAR for the Italian economy. Additionally, we evaluate non linear effects in the transmission of the shock to firms’ prices by using state dependent local projections framework (Ramey and Zubairy, 2018; Falck, Hoffmann and Hürtgen, 2021). Our contribution to the literature lies in examining how natural gas price shocks affect firm-level outcomes using microdata. We base our analysis on quarterly survey data from the Bank of Italy’s Survey on Inflation and Growth Expectations (SIGE).
@@
We draw from a vast literature about macroeconomic outcomes of energy related shocks. A large number of papers study the macroeconomic effects of oil shocks (Hamilton, 1983; Kilian, 2009; Caldara, Cavallo and Iacoviello, 2019; Conflitti and Luciani, 2019; Kilian and Zhou, 2022). Our work is closely related to the recent papers which study the effect of gas price and supply shocks, focusing on the recent inflationary surge in European economies. Alessandri and Gazzani (2025) identify a natural gas supply shock using daily news on European gas market as an instrument. Boeck and Zörner (2025) and López et al. (2025) estimate the pass through to inflation of gas price shocks. Adolfsen et al. (2024) use a BVAR framework to identify shocks driving the natural gas market in the EU. They document that the pass through is heterogeneous depending on the shock type. Güntner, Reif and Wolters (2024) study the effect of recent supply disruptions in natural gas supply in the German market. Casoli, Manera and Valenti (2024) study the interaction between oil/gas shocks and their effects on inflation in the Euro area. With respect to those authors, we focus on firm level expectations and price setting behavior.
Secondly, our work relates to the literature on uncertainty Bloom, Bond and Van Reenen (2007); Bloom (2009), especially with regard to that related to measuring uncertainty in agents’ expectations (Binder, 2017; Jurado, Ludvigson and Ng, 2015; Rossi and Sekhposyan, 2015; Manski, 2018) and to quantifying the effects of uncertainty on the macro economy (Bloom et al., 2018; Ascari and Haber, 2022; Georgarakos et al., 2024; Fasani et al., 2025).
The rest of the paper is organized as follows. In section 2 we describe the main features in the Survey of Inflation and Growth Expectaion (SIGE), which constitute our main data source of information for firm level microdata. We then describe the aggregate statistics obtained from the survey and that we use for conducting our empirical analysis, which is described in section 3. section 4 concludes.
@@ 2 Data: Survey on Inflation and Growth Expectations
To determine which factors are most likely to impact firms’ price-setting behavior, we examine their responses to the categorical survey questions. Firms are asked about both the direction (positive, negative) and the intensity (strong, medium, modest) with which a specific factor is likely to affect their prices in the next year. The factors that are considered are the prices of raw materials and intermediate inputs, their inflation expectations for the next year, the prices of other competing firms, the trend in labor cost and aggregate demand and changes in the situation related to financing conditions. By using these responses, we build diffusion indexes using the approach of Pinto, Sarte and Sharp (2020). Those indexes, shown in Figure 1 together with 95% confidence bands, capture the average perceived intensity and direction of each factor across firms and are informative about movements in the distribution of firms’ responses. [^2] In order to buid the indexes, each response is weighted by its intensity on a -3 to +3 scale, where higher absolute values indicate stronger perceived pressure.
Figure 1: Diffusion indexes on factors affecting firms’ future prices. The figure reports diffusion indexes computed on the categorical questions in the SIGE survey about the factors that can affect a firm’s prices in the next year, togheter with 95% confidence bands. Source: Survey of Inflation and Growth Expectation. Sample: 1999:Q4–2025:Q2.
The observed patterns suggest the presence of a cost-push narrative during the European energy crisis that followed the post-pandemic recovery: raw material and intermediate input costs are cited as primary pricing drivers starting in 2021, followed by labor costs and demand pressures. Notably, inflation expectations themselves become a significant pricing factor only around 2022, suggesting that as firms observed widespread price increases, expectations began feeding directly into their own pricing decisions. Given the prominence of input costs—particularly energy-related raw materials—as key drivers of pricing behavior during this period, this motivates our focus on examining the effects of natural gas price shocks on firm-level outcomes.
@@
In this paragraph, we report the first moment of firms’ inflation expectations during the next year, firms expectations about the percentage of change in the price they will charge in the next year and the percentage of change of the price they charged during the current year. Starting from the third quarter of 2012, an important innovation was introduced in the survey. Prior to this change, all participating firms were informed about the current level of inflation, which was reported in the survey questionnaire. Since 2012Q3, the sample size has been increased and firms have been randomly assigned to two groups: one receiving updated information about current inflation, and the other receiving no such information. Due to the lenght of the sample size, the aggregate statistics we report are those related to firms which have been updated about the current level of prices. It is then not surprising that the average annual inflation expectations reported in Figure 2 is closely tied to realized inflation. [^3] This stress the importance of lack of information and inattention as key sources of bias in inflation expectations.
Figure 2: Average firms’ annual inflation expectations. The figure reports the first moment of firms’ year-on-year inflation expectations, as well as the realized HICP. Source: Survey of Inflation and Growth Expectation. Sample: 1999:Q4–2025:Q2.
However, it is possible to observe a detachment from actual realizations during periods characterized by huge levels of economic distress. During the Leheman crisis of 2008 and the sovereign debt crisis firms tend to over predict inflation, while in the aftermarth of the Covid pandemic they massively under predicted it.
Figure 3: Average change in firms’ annual expected and realized prices. The figure reports the first moment of firms’ year-on-year expectations on their own prices, as well as the realized change in the price they charge and HICP. Source: Survey of Inflation and Growth Expectation. Sample: 1999:Q4–2025:Q2.
@@
To measure firms’ inflation uncertainty, we adopt the approach proposed by Binder (2017), which exploits the well-documented tendency of survey respondents to provide round-number forecasts when they face greater uncertainty. This method yields an index that captures the proportion of likely uncertain firms in each survey wave. Since firms are informed about current price levels at the time of the interview, our index isolates doubts about future inflation, abstracting from any confusion about the current state of prices. Figure 4 reports inflation uncertainty index between the last quarter of 1999 and the second quarter of 2025.
@@
Figure 4: Inflation Uncertainty Index. The figure reports the inflation uncertainty index for firms computed by adopting the framework of Binder (2017). Vertical red bars higlight moments of high uncertainty. Source: Survey of Inflation and Growth Expectation. Sample: 1999:Q4–2025:Q2.
@@
Inflation uncertainty among informed firms is notably lower than that typically observed for consumers (Binder, 2017). During the early sample period, uncertainty remains low, reflecting stable price dynamics. The first significant spike occurs in 2008Q3 during the Global Financial Crisis (GFC), followed by elevated levels during the sovereign debt crisis. Subsequent peaks coincide with the Russian invasion of Crimea (2014Q3), Greece’s IMF default (2015Q2), and the peak in the Italian banking crisis (2016Q3). The fall of the Conte II government in 2020Q4 triggers a sharp increase in uncertainty, likely reflecting concerns about Italy’s access to European recovery funds during the COVID-19 pandemic. The most dramatic increase in uncertainty emerges with Russia’s invasion of Ukraine in 2022Q1, driven by severe energy supply disruptions and their inflationary consequences. Uncertainty peaks again in 2022Q4, as Italian inflation reaches its highest level since the introduction of the Euro, reflecting widespread concerns about the persistence and breadth of price pressures.

### v3

# Gas supply shocks, Uncertainty and Price setting: Evidences from Italian Firms

@@ Abstract
This paper examines how natural gas supply shocks affect Italian firms’ pricing decisions and inflation expectations using quarterly survey data from the Bank of Italy’s Survey on Inflation and Growth Expectations (SIGE). We identify natural gas supply shocks through an external IV-VAR approach exploiting likely unexpected news about interruption to gas supplies to Europe. Our findings show that although gas supply shocks do not have huge effects on gas quantity and only modest effect on gas inventories, they are quickly transmitted to spot electricity prices with persistent effects. We then estimate a proxy internalizing BVAR incorporating firm-level variables from SIGE, documenting that gas supply shocks raise firms’ current and expected prices as well as inflation uncertainty. Finally, we uncover substantial nonlinearities using state-dependent local projections: under high inflation uncertainty, firms successfully pass cost increases on to consumers, sustaining elevated prices; under low uncertainty, recessionary effects dominate, leading firms to cut prices below baseline.
Keywords: Gas supply shocks, Firms’ Price Setting, Inflation Uncertainty, Firms Expectations, Local Projections, Bayesian Vector Autoregression, Instrumental Variables
@@ 1 Introduction
The period that follows the Covid-19 pandemic has seen an unprecedented rise in inflation in European economies. The post pandemic recovery has been characterized by strong political tensions, dominated by events such as Russian invasion of Ukraine in the first part of 2022 which led to massive disruptions in natural gas supplies to European countries, raising the attention on inflationary role of supply shocks affecting this important commodity. Due to its limited domestic energy resources, Italy is a major importer and remains highly dependent on natural gas as a primary source of power, making it particularly sensitive to the supply disruptions that followed this dramatic geopolitical turning point. Italian firms have been particularly exposed to the unprecedented increase in the price of natural gas, which had a substantial impact on the price of electricity in the country. We show that, unsurprisingly, the post-pandemic recovery period has been characterized by an unprecedented increase in firms’ uncertainty about the future level of prices.
Unlike the extensive literature on oil supply shocks, research examining the role of natural gas supply disruptions has emerged only recently. In this paper, we aim at identifying a shock to the supply of natural gas and at evaluating the impact of this shock on firms price setting behavior. We identify the shock following an external IV-VAR approach following kanzig2021macroeconomic. We construct the instrument by exploiting episodes that carried likely exogenous news about supply distruptions affecting Europe, and in particular Italy. Our identification strategy is hence close in the spirit to the one followed by alessandri2025natural, which instead adopt an internal instrument and consider a larger number of surprises. Then, we evaluate the effect of this shock on firms’ pricing behavior and expectations using the generalized impulse response functions from a proxy internalizing VAR, using the shock identified in the first VAR as an external instrument. Additionally, we evaluate non linear effects in the transmission of the shock to firms’ prices by using state dependent local projections framework (ramey2018government; F2021). Our contribution to the literature lies in examining how natural gas supply shocks affect firm-level outcomes and the pass-through using firm level micro data. We base our analysis on quarterly data from the Bank of Italy’s Survey on Inflation and Growth Expectations (SIGE).
@@
We draw from a vast literature about macroeconomic outcomes of energy related shocks. A large number of papers study the macroeconomic effects of oil shocks (hamilton1983oil; kilian2009not; caldara2019oil; conflitti2019oil; kilian2022oil). Our work is closely related to the recent papers which study the effect of gas supply shocks, focusing on the recent inflationary surge in European economies. alessandri2025natural identify a natural gas supply shock using daily news on European gas market as an instrument. boeck2025natural and lopez2025pass estimate the pass through to inflation of gas price shocks. adolfsen2024gas use a BVAR framework to identify shocks driving the natural gas market in the EU. They document that the pass through is heterogeneous depending on the shock type. guntner2024sudden study the effect of recent supply disruptions in natural gas supply in the German market. casoli2024energy study the interaction between oil/gas shocks and their effects on inflation in the Euro area. With respect to those authors, we focus on firm level expectations and price setting behavior.
Secondly, our work relates to the literature on uncertainty bloom2007uncertainty; Bloom2009, especially with regard to that related to measuring uncertainty in agents’ expectations (binder2017measuring; Jurado2015; RossiTatevik; Manski2018) and to quantifying the effects of uncertainty on the macro economy (Bloom2018; ascari2022non; GGCK2024; fasani2025belief).
The rest of the paper is organized as follows. In section 2 we describe the main features in the Survey of Inflation and Growth Expectation (SIGE), which constitute our main data source of information for firm level micro data. We then describe the aggregate statistics obtained from the survey and that we use for conducting our empirical analysis, which is described in section 3. section 4 concludes.
@@ 2 Data: Survey on Inflation and Growth Expectations
To determine which factors are most likely to impact firms’ price-setting behavior, we examine their responses to the categorical survey questions. Firms are asked about both the direction (positive, negative) and the intensity (strong, medium, modest) with which a specific factor is likely to affect their prices in the next year. The factors that are considered are the prices of raw materials and intermediate inputs, their inflation expectations for the next year, the prices of other competing firms, the trend in labor cost and aggregate demand and changes in the situation related to financing conditions. By using these responses, we build diffusion indexes using the approach of PintoSarteSharp2020. Those indexes, shown in Figure 1 together with 95% confidence bands, capture the average perceived intensity and direction of each factor across firms and are informative about movements in the distribution of firms’ responses. [^2] In order to build the indexes, each response is weighted by its intensity on a -3 to +3 scale, where higher absolute values indicate stronger perceived pressure.
Figure 1: Diffusion indexes on factors affecting firms’ future prices. The figure reports diffusion indexes computed on the categorical questions in the SIGE survey about the factors that can affect a firm’s prices in the next year, together with 95% confidence bands. Source: Survey of Inflation and Growth Expectation. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 1999:Q4–2025:Q2.
The observed patterns suggest the presence of a cost-push narrative during the European energy crisis that followed the post-pandemic recovery: raw material and intermediate input costs are cited as primary pricing drivers starting in 2021, followed by labor costs and demand pressures. Notably, inflation expectations themselves become a significant pricing factor only around 2022, suggesting that as firms observed widespread price increases, expectations began feeding directly into their own pricing decisions. Given the prominence of input costs—particularly energy-related raw materials—as key drivers of pricing behavior during this period, this motivates our focus on examining the effects of natural gas supply shocks on firm-level outcomes.
@@
In this paragraph, we report firms’ one-year-ahead inflation expectations, their expected price changes over the same horizon, and their actual price changes during the current year. Starting from the third quarter of 2012, an important innovation was introduced in the survey. Prior to this change, all participating firms were informed about the current level of inflation, which was reported in the survey questionnaire. Since 2012Q3, the sample size has been increased and firms have been randomly assigned to two groups: one receiving updated information about current inflation, and the other receiving no such information. Due to the length of the sample size, the aggregate statistics we report are those related to firms which have been updated about the current level of prices. It is then not surprising that the average annual inflation expectations reported in Figure 2 is closely tied to realized inflation. [^3] This stress the importance of informational frictions and inattention as key sources of bias in inflation expectations.
Figure 2: Average firms’ annual inflation expectations. The figure reports the first moment of firms’ year-on-year inflation expectations, as well as the realized HICP. Source: Survey of Inflation and Growth Expectation. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT. Sample: 1999:Q4–2025:Q2.
However, it is possible to observe a detachment from actual realizations during periods characterized by huge levels of economic distress. During the Lehman crisis of 2008 and the sovereign debt crisis firms tend to over predict inflation, while in the aftermath of the COVID pandemic they massively under predicted it.
Figure 3: Average change in firms’ annual expected and realized prices. The figure reports the first moment of firms’ year-on-year expectations on their own prices, as well as the annual realized change in the price they charge and realized HICP. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT. Sample: 1999:Q4–2025:Q2.
@@
To measure firms’ inflation uncertainty, we adopt the approach proposed by binder2017measuring, which exploits the well-documented tendency of survey respondents to provide round-number forecasts when they face greater uncertainty. This method yields an index that captures the proportion of likely uncertain respondents in each survey wave. Differently from her case, our respondents can provide decimal answers. Therefore, we need to adapt her framework to our setting by adjusting the likelihood function accordingly. We assume that there are two type of respondents. TypehType-h respondents are characterized by an high degree of uncertainty in their subjective probability distribution, while typeltype-l respondents by a low degree of uncertainty. TypehType-h respondents are assumed to give point estimates for their expected level of prices coming from the set $S_{h}=\{0,5,10,15,20,25\}$, while low uncertainty firms are allowed to answer with decimal numbers on the support $\mathbb{R}^{+}$. If an answer $R_{it}\notin S_{h}$ we can surely classify it as typeltype-l. Conversely, if $R_{it}\in S_{h}$ we need to estimate the probability that respondent $i$ is uncertain. We assume that the cross sectional density is the mixture of two distributions. For typehtype-h respondents, coming from $S_{h}$ the probability mass function is obtained by integrating the latent density over rounding bins: $\phi_{h}(s)=P(R_{i,t}=s|type-h)=\int_{s-\delta/2}^{s+\delta/2}p_{h}(x)dx$, with $j\in S_{h}$ and $\delta$ being the bin width. [^4] We set $\delta=0.1$. For typeltype-l respondents the probability density is simply $f_{l}(x)=p_{l}(y)$ with $y\in\mathbb{R}^{+}$ and $\phi_{l}(s)=\int_{s-\delta/2}^{s+\delta/2}p_{l}(x)dx$ for rounded observations. The cross sectional log likelihood will then be given by:
$l(\theta_{t})=\sum_{s=S_{h}}n_{h}(s)log[(\lambda_{t}\phi_{h}(s)+(1-\lambda_{t})\phi_{l}(s)]+\sum_{i=1}^{N_{c}}log[(1-\lambda_{t})f_{l}(y_{c,i};\mu_{l},\sigma_{l})]$ (1)
where $f_{l}$ is assumed to be normal, $n_{h}$ is the number of firms which responses belong to $S_{h}$, $N_{c}$ is the number of non rounded responses and $\theta=(\phi_{h},\lambda_{t},\mu_{l},\sigma_{l})$ is a vector of parameters . The mixture weight $\lambda_{t}$ is the portion of uncertain respondents.555binder2017measuring includes non respondents and measures the proportion of uncertain households as $U_{t}=\frac{\lambda_{t}N_{t}+\gamma N_{t}^{DK}}{N_{t}+\gamma N_{t}^{DK}}$ where $N_{t}$ is the total number of respondents, $N_{t}^{DK}$ is the number of respondents which answer is ”don’t know” and $\gamma$ is a parameter which measures their weight. As in the SIGE survey the question about expected inflation is mandatory, we do not have DKtypeDK-type respondents in our sample and hence we measure the proportion of uncertain firms by simply taking $\lambda_{t}$. The vector of parameters $\theta_{t}$ is estimated via maximum likelihood. Figure 4 reports the inflation uncertainty index between the last quarter of 1999 and the second quarter of 2025. Notably, since firms are informed about current price levels at the time of the interview, our index isolates doubts about future inflation, abstracting from any confusion about the current state of prices.
@@
Figure 4: Inflation Uncertainty Index. The figure reports the inflation uncertainty index for firms computed by adopting the framework of binder2017measuring. Vertical red bars highlight moments of high uncertainty. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 1999:Q4–2025:Q2.
@@
Inflation uncertainty among informed firms is notably lower than that typically observed for consumers (binder2017measuring). During the early sample period, uncertainty remains low, reflecting stable price dynamics. The first major spike occurs in 2008Q3 during the Global Financial Crisis, followed by persistently high levels throughout the European sovereign debt crisis. Subsequent surges are observed around the Russian invasion of Crimea (2014Q3), Greece’s IMF default (2015Q2), and during the Italian banking crisis (2016Q3). The fall of the Conte II government in 2020Q4 coincides with a sharp rise in uncertainty, likely reflecting concerns over Italy’s access to European recovery funds amid the COVID-19 pandemic. A major increase appears in 2022Q1 with Russia’s invasion of Ukraine, driven by severe energy supply disruptions and their inflationary effects. Uncertainty rises again in 2022Q4, as Italian inflation reaches its highest level since the adoption of the euro, underscoring widespread concerns about the persistence and spread of price pressures.